the measure of an angle is 58 degrees more than the measure of its supplement. Find the measure of its supplement.
Answer by Paul(988)

58+x+x=180
2x=180-58
2x=122
x=61
61+58=119
Hence, the two angles are 119 and 61. The supplement is 61 degrees.